40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 5 18:57:10 2011, Jamie, Update, SUS, More normalization of all sus controllers 
    Reply  Tue Jul 5 21:26:51 2011, rana, Update, SUS, More normalization of all sus controllers Screenshot-1.png
       Reply  Wed Jul 6 02:12:36 2011, kiwamu, Update, SUS, Re : More normalization of all sus controllers 
          Reply  Wed Jul 6 10:35:35 2011, Jamie, Update, SUS, Re : More normalization of all sus controllers 
             Reply  Wed Jul 6 15:32:32 2011, Jamie, Update, SUS, Re : More normalization of all sus controllers 
       Reply  Wed Jul 6 11:45:20 2011, Jamie, Update, SUS, More normalization of all sus controllers 
Message ID: 4943     Entry time: Wed Jul 6 02:12:36 2011     In reply to: 4942     Reply to this: 4944
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: Re : More normalization of all sus controllers 

[Jenne / Rana/ Kiwamu]

 We found the 30 Hz high pass filters had lower gain than what they used to be at low frequcnies.

So we increased the gain of the high pass filters called '30:0.0'  by a factor of 10 to have the same gain as before.

Now all the suspension shows some kind of damping. Needs more optimizations, for example Q-adjustments for all the suspensions...

Quote from #4942

This is getting closer, but with the whitening left OFF and the cts2um filter also OFF, none of the suspensions are working correctly.

 

ELOG V3.1.3-